at the fair you win $100 if you ring the loop on the hook. if not you win $0 the probability is very low at 2% chance determine how much you will expect to win in one try?
1
Expert's answer
2020-04-27T17:56:28-0400

x=winning price in one try

x=100 corresponding probability= 0.02

x=0 corresponding probability= 0.98

"E[x]=\sum x.p(x)=100*0.02+0*0.98=2"


Therefore expected to win in one try is $2